OpenStudy (vinnybrown1997):

An investment account earns 4% per year compounded annually. If the initial investment was $4,000.00, how much is in the account after 3 years? Round your answer to the nearest dollar

OpenStudy (ltrout):

Use the formula \[A = P(1+r/n)^{nt}\] Where: P = initial amount invested - 4000 R is the rate 0.04 N is the number of times it is compounded for year (1). T is the time, 3 So the equation would be: \[A = 4000(1.04)^{1(3)}\] Can you solve from there?